OpenStudy (anonymous):

1) A 30-kg block sits at rest on a horizontal surface. The coefficient of static friction μs = 0.60. The coefficient of kinetic friction μk = 0.5. A force P of 250 N is applied in the +x direction. A) The block remains at rest. B) The block moves and continues to move at constant velocity. C) The block accelerates to the right. D) The block does not move until the force is increased to greater than 500 N. E) No conclusions can be drawn concerning the movement of the block from the information given.

OpenStudy (anonymous):

How do I do this?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

the limiting force to move the block is μmg = 0.6 x 30 x 10 = 180 since applied force is greater than that, block will move in direction of applied force... when it moves the force equation is F- f = ma F is the applied force, f the kinetic friction...hence it accelerates 250 - (0.5 x 30 x 10) = 30a a=10/3